OpenStudy (anonymous):

Solve 1/F=1/m+1/p for m

OpenStudy (bahrom7893):

multiply everything out by m: m/F = 1 + m/p

OpenStudy (bahrom7893):

Move all the terms with m to one side: m/F - m/p = 1

OpenStudy (bahrom7893):

pull out an m: m(1/F - 1/p) = 1

OpenStudy (bahrom7893):

Divide both sides by 1/F-1/p: m = 1/(1/F-1/p). Now just do some simplification to get your answer.